# ID-Guard: A Universal Framework for Combating Facial Manipulation via Breaking Identification

### v1

@@ III-A2 Cross-Model Universal Perturbations
where $\mathcal{S}_{\lambda}=\left\{\lambda_{1},\lambda_{2},\dots,\lambda_{N}\right\}\in\mathbb{R}$ is a set of weights got based on the robustness of each target model. Solving the effective $\mathcal{S}_{\lambda}$ is one of the key tasks of this work so that the generated adversarial perturbations have a balanced defense performance against different facial manipulations. This part will be introduced in detail in Section III-C.
Figure 4: Illustration of the proposed ID-guard framework. The perturbation generator takes a natural image $x$ as input and requires only one forward propagation to generate a cross-model adversarial perturbation dedicated to the input face that can be used to defend against multiple facial manipulations. In the training phase, ID-guard consists of three modules, including the identity destruction module, the dynamic weighting strategy, and the gradient prior perturbation strategy. The notation descriptions are shown in the upper right corner for reference.
@@ III-B Identity Destruction Module
As mentioned above, the design of the distance metric $\mathcal{D}$ is very critical. Traditional methods [10, 11, 39, 12] generally choose Mean Squared Error (MSE) as a proxy for distance measurement. However, this approach causes non-directional distortion of forged images and fails to ensure that the identity of the distorted face cannot be recognized. As an improvement, we introduce the Identity Destruction Module (IDM). As shown in Fig. 4, the IDM consists of three sub-modules, which will be introduced separately next.
III-B1 Mask Loss
First, we consider using face masks to limit the regions of image distortion by adversarial perturbations. The designed mask is two-fold: 1) The binary mask is used to restrict image distortion to areas of facial components including the eyes, nose, mouth, and eyebrows, which are proven to play an important role in identity recognition by the human eye [40, 41, 42]; 2) The heatmap mask weights the face distortion loss at the pixel level, making the perturbation pay more attention to the important feature areas of the face. In this work, the heatmap of each image is obtained by solving Grad-SAM [43] on VGGFace [44]. This design will also facilitate distorted images against commercial facial recognition systems. Hence, for the facial manipulation model $\mathcal{M}_{k}$, the mask loss can be formulated as:
@@ III-B1 Mask Loss
where $m^{bin}$ and $m^{hm}$ denote the binary mask and heatmap mask of the natural image $x$, respectively. Note that these masks are only computed during the perturbation generator training stage to constrain the distortion region and are not used in the inference process. $\odot$ indicates the element-wise multiplication.
III-B2 Identity Loss
In addition to pixel-level constraints, further consideration is given to maximizing the identity similarity of the manipulation results on the original and adversarial images. We achieve this by maximizing the cosine distance between their identity embeddings:
@@ III-B2 Identity Loss
III-B3 Feature Loss
Learning from [17] that the process of attribute or identity embedding for end-to-end facial manipulation varies, but the feature extraction process is similar. In addition, feature-level perturbations can retain their effective components in network transmission to a greater extent [16]. Therefore, a feature loss that enables the generator to focus on destroying feature-level faces is incorporated into the IDM to improve the effectiveness and transferability of produced perturbations:
@@ III-B3 Feature Loss
In summary, given a pre-trained facial manipulation model $\mathcal{M}_{k}\in\mathcal{S}_{\mathcal{M}}$, the attacking loss against it can be formulated as:
$\mathcal{L}_{k}=\mathcal{L}_{k}^{mask\_bin}+\mathcal{L}_{k}^{mask\_hm}+\mathcal{L}_{k}^{ID}+\mathcal{L}_{k}^{feat}$ (8)
The loss function for training the perturbation generator $\mathcal{G}$ is a linear combination of attacking losses of facial manipulation models:
$\mathcal{L}=\lambda_{1}\cdot\mathcal{L}_{1}+\lambda_{2}\cdot\mathcal{L}_{2}+\dots+\lambda_{N}\cdot\mathcal{L}_{N}$ (9)
@@ III-C Dynamic Weight Strategy
We regard the attack against different facial manipulation models in Eq. (9) as a multi-task learning problem to solve the weight set $\mathcal{S}_{\lambda}$. As shown in Fig 4, the proposed dynamic weighting strategy is two-fold. First, a classical but effective multi-task learning strategy, i.e., MGDA [28], is integrated into the proposed ID-guard framework in order to dynamically adjust the weights to optimize the gradients of multiple tasks during the generator training phase. Specifically, we follow [27]  [^4] https://github.com/isl-org/MultiObjectiveOptimization to solve this multi-objective optimization problem and use it as a baseline dynamic weighting strategy (noted as S-I).
On the other hand, we introduce a strategy that combines prior weights and dynamic weights (noted as S-II). For the $\lambda_{k}$ of the attack loss in Eq. (9), refine it to
@@
where $t$ indexes the iteration steps, $\alpha_{k}=10^{n}\,(n\in\mathbb{Z})$ is a prior order of magnitude weight for the attack loss, which is determined based on the adversarial robustness of the corresponding facial manipulation model. Here, a heuristic is used to quantify the adversarial robustness: on a mini-batch of images, the equivalent setting PGD [35] is used to derive a gradient-based adversarial attack against facial manipulation models, and the $L_{2}$ distance of the attack is calculated. $\beta_{k}^{t}$ is dynamically adjusted during the training stage. Inspired by [31], we would like to make the harder-to-learn task, i.e., the harder-to-attack facial manipulation model, have higher weights during training. Hence, $\beta_{k}^{t}$ is solved as:
$\beta_{k}^{t}=-(1-\mathcal{K}_{k}^{t})\log\mathcal{K}_{k}^{t}$ (11)
where $\mathcal{K}_{k}$ represents the KPI for $\mathcal{M}_{k}$, which is an indicator of attack performance. KPI is inversely proportional to the training difficulty of the task, that is, the higher the KPI, the easier the task is to learn. Here we choose $L_{2}$ distance as the proxy of KPI. Intuitively, tasks with high KPIs are easier to learn and therefore become less weighted; conversely, tasks that are difficult to learn become more weighted. This combined design improves the stability of dynamic weights to produce adversarial perturbations with more balanced performance.
@@ III-D Gradient Prior Perturbation
where $\mathrm{cat}(\cdot)$ denotes channel-wise concatenation, i.e., $\mathrm{cat}(x,\delta_{p})\in\mathbb{R}^{6\times H\times W}$. The intuition behind this design is that the gradient prior perturbation can provide the generator with rich prior gradient and structural perturbation information, thereby promoting more stable training and more efficient perturbation generation.
@@ IV-A1 Datasets
In our experiments, the CelebAMask-HQ [47] dataset is selected for training the perturbation generator. It consists of more than 30,000 face images, where each image carries semantic masks for 19 facial component categories. These fine-grained mask labels can provide support for computing the binary mask loss during the training stage. To adequately evaluate the performance as well as the generalizability of our method and the competing algorithms, we test them on three datasets including CelebAMask-HQ [47], LFW [48] and FFHQ [49].
@@ IV-A2 Target Models
IV-A3 Baselines
To demonstrate the superiority of the proposed method in face identity protection and cross-model universal performance, four advanced proactive defense methods including Disrupting [10], PG [39], CMUA [12] and IAP [15] are selected as competing algorithms. Disrupting [10] disrupts facial manipulation by iteratively solving gradient-based adversarial perturbations on the target model. PG [39] achieves adversarial perturbation generation in gray-box scenarios by attacking a surrogate model. CMUA [12] is a baseline of universal defense against multiple models. IAP [15] designs an information-containing adversarial perturbation, but we only implement its proactive distortion without involving the embedding and extraction of information to provide a fair comparison.
IV-A4 Metrics
where $(i,j)$ is the coordinate of pixels and $Face_{i,j}$ is the binary facial mask of the image. The pixel value of its face area is 1, otherwise it is 0. The binary facial mask is calculated by Dlib  [^5] https://pypi.org/project/dlib. Additionally, we evaluate the identity similarity (noted as ID sim. in Tables) computed by Arcface [45] between the forged and distorted outputs. Defense success rates are also considered. Previous works [10, 12] have generally determined the success of a defense by whether the $L_{2}$ distance is greater than 0.05, but this is incomplete in the task of preventing face stigmatization. As shown in Fig. 5, the distorted output in Fig. 5 (c) reports a successful defense at the $L_{2}^{face}$ distance, but it appears that it merely “blackens” the face without destroying the individual’s identity. Therefore, we propose that both $L_{2}^{face}$ distance greater than 0.05 and identity similarity less than 0.4 be satisfied to indicate successful defense, which is a more challenging evaluation. The contrast between (d) and (e) in Fig. 5 shows the necessity of considering both restrictions simultaneously.
IV-A5 Implementation Details
@@
Ours (S-II) 0.536 0.002 1.000 0.329 0.107 0.970 0.502 0.025 1.000 0.345 0.030 0.995 0.183 0.081 0.964
@@ IV-B Comparison with Baselines
Table I summarizes the quantitative comparison of the proposed ID-guard with competitive algorithms. Our method is reported separately under two strategies, as presented in III-C. The Disrupting [10] produces gradient-based adversarial perturbations against StarGAN [21], which makes it overfit in disrupting this model at the cost of cross-model performance. The perturbation optimization of CMUA [12] and PG [39] is unconstrained and thus has limited destruction to the identity semantics. As shown in Fig. 6, the person’s identity in a distorted image under the CMUA can be still recognized. Attributed to the feature correlation measurement loss employed, IAP [15] improves the performance to destroy identification to some limit extent. In comparison, our method significantly destroys the identity semantics of images and effectively prevents the stigmatization of faces. Furthermore, the baseline methods equally weight the attack loss of different facial manipulations, leading to perturbations biased towards vulnerable models. Due to the introduction of the dynamic parameter strategy, we achieve balanced performance against various facial manipulations. For the most robust model against attack, HiSD [25], ID-guard improves the defense success rate by 50% compared with the state-of-the-art method.
On the three selected datasets, ID-guard under both strategies demonstrated superior performance. For strategy I, the MGDA algorithm is employed to adjust the weights, which has the advantage of completely eliminating the need for human intervention and prior knowledge during the training process. However, this non-intervention makes ID-guard based on strategy I exhibit an extreme effect to some extent: it performs better on the more vulnerable model (e.g., AGGAN [22]) and the more robust model (e.g., HiSD [25]), but does not provide significant improvement on models in the middle (e.g., RelGAN [24]). Moreover, additional backpropagation computation is required at each iteration when implementing MGDA, which increases the training overhead. In contrast, Strategy II-based ID-guard only additionally computes a set of KPI values, which has a minimal impact on training overhead. Strategy II maintains a set of prior parameters for balancing each model, thereby further balancing performance. The disadvantage of it is the extra work of determining the prior parameter set.
Figure 6: Visual examples of disruption to different facial manipulations. Compared with CMUA [12], which still retains a large amount of identity semantics in the distorted image, our method makes the face in the image completely unrecognizable.

### v2

@@ III-A3 ID-Guard Framework
where $\mathcal{S}_{\lambda}=\left\{\lambda_{1},\lambda_{2},\dots,\lambda_{N}\right\}$ represents a set of weights used to balance the adversarial loss for different models during training. Determining these weights is one of our key tasks. To overcome the adversarial gradient discrepancies caused by significant structural differences among facial manipulation models, we propose a dynamic weight strategy to achieve balanced cross-model defense performance, which will be introduced in Section III-C. Additionally, to address the challenge of facial stigmatization caused by non-specific constraint distortions, we propose an Identity Destruction Module (IDM) to compute the distance metric $\mathcal{D}$, which will be detailed in Section III-B. Moreover, to incorporate structural adversarial gradient priors during the training of $\mathcal{G}$, we introduce a Gradient Prior Strategy, which is discussed in Section III-D.
Figure 4: Illustration of the proposed ID-Guard framework. The perturbation generator takes a natural image $x$ as input and requires only one forward propagation to generate a cross-model adversarial perturbation dedicated to the input face that can be used to defend against multiple facial manipulations. In the training phase, ID-Guard consists of three modules, including the Identity Destruction Module, the dynamic weight strategy, and the gradient prior perturbation strategy. The notation descriptions are shown in the upper right corner for reference.
@@ III-B Identity Destruction Module
To clarify the algorithmic design, we first introduce the Identity Destruction Module (IDM). Based on this module, we then provide a complete definition of the training loss function for the perturbation generator $\mathcal{G}$. As shown in Fig. 4, the IDM consists of three sub-modules, which will be introduced separately next.
III-B1 Mask Constrained Loss
First, we consider using face masks to limit the regions of image distortion by adversarial perturbations. The designed mask is two-fold: 1) The binary mask is used to restrict image distortion to areas of facial components including the eyes, nose, mouth, and eyebrows, which are proven to play an important role in identity recognition by human eyes [42, 40, 41]; 2) The heatmap mask weights the face distortion loss at the pixel level, making the perturbation pay more attention to the important feature areas of the face. In this work, the heatmap of each image is obtained by solving Grad-cam [45] on VGGFace [44]. This design will also facilitate distorted images against commercial facial recognition systems. Hence, for the facial manipulation model $\mathcal{M}_{k}$, the mask loss can be formulated as:
@@ III-B1 Mask Constrained Loss
where $m^{bin}$ and $m^{hm}$ denote the binary mask and heatmap mask of the original image $x$, respectively. Note that these masks are only computed during the perturbation generator training stage to constrain the distortion region and are not used in the inference process. $\odot$ indicates the element-wise multiplication.
III-B2 Identity Consistency Loss
In addition to pixel-level constraints, we also consider maximizing the identity discrepancy between the forgery outputs of the facial manipulation model $\mathcal{M}_{k}$ for the original image and the adversarial image, i.e., maximizing the discrepancy between the original forged image $\mathcal{M}_{k}(x)$ and the disrupted forged image $\mathcal{M}_{k}(x_{adv})$. This is achieved by maximizing the cosine distance between their identity embeddings:
@@ III-B2 Identity Consistency Loss
III-B3 Feature Confusion Loss
It is crucial to exploit the commonalities between different facial manipulation models during perturbation generation. Learning from [25] that while the attribute embedding processes in end-to-end facial manipulation models vary, their feature extraction processes share similarities. In addition, feature-level perturbations can retain their effective components in network transmission to a greater extent [15]. Therefore, a feature confusion loss that enables the generator to focus on destroying feature-level faces is incorporated into the IDM to improve the effectiveness and transferability of produced perturbations:
@@ III-B3 Feature Confusion Loss
In summary, given a pre-trained facial manipulation model $\mathcal{M}_{k}\in\mathcal{S}_{\mathcal{M}}$, the adversarial loss against it can be formulated as:
$\mathcal{L}_{k}^{adv}=\mathcal{L}_{k}^{mask\_bin}+\mathcal{L}_{k}^{mask\_hm}+\mathcal{L}_{k}^{ID}+\mathcal{L}_{k}^{feat}$ (9)
The loss function for training the perturbation generator $\mathcal{G}$ is a linear combination of adversarial losses of facial manipulation models:
$\mathcal{L}^{adv}=\lambda_{1}\cdot\mathcal{L}_{1}^{adv}+\lambda_{2}\cdot\mathcal{L}_{2}^{adv}+\dots+\lambda_{N}\cdot\mathcal{L}_{N}^{adv}$ (10)
@@ III-C Dynamic Weight Strategy
We regard the attacks against different facial manipulation models in Eq. (10) as a multi-task learning problem to optimize the weight set $\mathcal{S}_{\lambda}$, ensuring that the generated perturbations achieve more balanced cross-model adversarial effectiveness. As shown in Fig. 4, the proposed dynamic weight strategy is two-fold.
III-C1 MGDA-based
First, we integrate the Multiple Gradient Descent Algorithm (MGDA) [30] into the proposed ID-Guard. The goal of MGDA is to find a set of weights $\lambda_{k}$ that balances the trade-offs between tasks while respecting the constraints of Pareto optimality. It can be formulated as the following quadratic programming problem:
$\begin{split}&\min_{\lambda\geq 0}\quad\left\|\sum_{k=1}^{N}\lambda_{k}\mathcal{L}_{k}^{adv}\right\|^{2},\\ &s.t.\quad\sum_{k=1}^{N}\lambda_{i}=1\end{split}$ (11)
In this paper, we employ the Frank-Wolfe method [29] to solve the optimization problem. The core idea is to combine gradients from multiple tasks into a suitable descent direction that minimizes the losses across all tasks. Specifically, we follow the implementation outlined in [29]  [^4] https://github.com/isl-org/MultiObjectiveOptimization. This dynamic weight strategy, based on MGDA, is used as a baseline, referred to as Strategy I (S-I).
III-C2 KPI-based
Additionally, a dynamic weight strategy based on Key Performance Indicators (KPI) is introduced, denoted as Strategy II (S-II). Each dynamic weight $\lambda_{k}$ is refined into a prior weight and a learnable weight, as follows:
@@ III-C2 KPI-based
where $t$ indexes the iteration steps, $\alpha_{k}=10^{n}\,(n\in\mathbb{Z})$ is the prior weight representing the magnitude of the attack loss, which reflects the adversarial robustness of different forged models, and $\beta_{k}^{t}$ is the learnable weight updated adaptively based on defense performance during each iteration.
To set the prior weight $\alpha_{k}$ reasonably, a small-sample heuristic method is used to quantify the prior adversarial robustness of different facial manipulation models. Specifically, a small batch of face images is randomly selected, and slight attacks are applied to each model using PGD [37] with the same settings. The magnitude of the $L_{2}$ distance is used to determine the value of the prior weight. This approach provides a reasonable initialization of weights in the early stages of training, balancing the contribution of adversarial losses across different models.
For the learnable weight $\beta_{k}^{t}$, inspired by the study [34], we assign higher loss weights to more difficult-to-learn tasks, i.e., facial manipulation models that are harder to disrupt during training. The $\beta_{k}^{t}$ is computed as follows:
$\begin{split}&\beta_{k}^{t+1}=-(1-\mathcal{K}_{k}^{t})\log\mathcal{K}_{k}^{t},\\ &\beta_{k}^{0}=1\end{split}$ (13)
where $\mathcal{K}_{k}^{t}$ is the KPI representing the attack difficulty of model $\mathcal{M}_{k}$ at iteration $t$. A higher KPI indicates that the model is easier to attack, so a smaller weight $\beta_{k}^{t}$ is assigned, while a lower KPI $\mathcal{K}_{k}$ results in a larger weight to enhance the adversarial impact of the generated perturbation against the model $\mathcal{M}_{k}$. Thus, the KPI value is inversely proportional to the adversarial loss weight. A negative correlation function of $\mathcal{K}_{k}^{t}$ is used to calculate the $\beta_{k}^{t}$. For convenience, we use the proposed $\mathcal{L}_{k}^{mask\_hm}$ to compute the distance between the real and perturbed fake results as a proxy for KPI, with its range truncated to $(0,1]$ to ensure the monotonicity of the function. Unlike the MGDA-based dynamic weight strategy, which requires additional computation in each iteration, the KPI-based strategy directly uses the loss function value calculated in the previous iteration as the KPI to obtain the learnable weight for this iteration. This combination design is more efficient and enhances the stability of the dynamic weight.
@@ III-D Gradient Prior Perturbation
where $\mathrm{cat}(\cdot)$ denotes channel-wise concatenation, i.e., $\mathrm{cat}(x,\delta_{p})\in\mathbb{R}^{6\times H\times W}$. Both parts are calculated using the adversarial loss defined in Eq.10. The intuition behind this design is that the gradient prior perturbation provides the generator with rich information about the prior gradient and perturbation structure, thereby promoting more stable training and more efficient perturbation generation. In particular, the optimized prior adversarial perturbation $\delta_{p}$ (noted as P-Pert) can be viewed as a cross-model universal adversarial perturbation that can protect multiple images from multiple facial manipulation models. The complete training process is given in Algorithm 1.
input : Original dataset $S_{ori}$, facial manipulation model set $\mathcal{S}_{\mathcal{M}}$, max iteration $T$.
output : Optimized $\mathcal{G}*$ and $\delta_{p}*$.
1
2// Generating initial prior adversarial perturbations $\delta_{p}^{0}$;
3 $\delta_{p}^{0}\leftarrow 0$;
4 $\hat{S}_{ori}\leftarrow SelectBatch(S_{ori})$;
5 foreach $x^{i}\in\hat{S}_{ori}$ do
6       // Implementing the PGD;
7       $\delta^{i}\leftarrow PGD(x^{i})$;
8       $\delta_{p}^{0}\leftarrow\delta_{p}^{0}+\delta^{i}$ ;
9
10 end foreach
11
[…]
13
[…]
15 foreach $\lambda_{k}\in\mathcal{S}_{\lambda}^{0}$ do
16       $\lambda_{k}^{0}\leftarrow 1$;
17
18 end foreach
19// Training the Perturbation Generator;
20 for $j\leftarrow 1$ to $T$ do
21       $x_{b}\leftarrow SelectBatch(S_{ori})$;
22       // Calculating the prior weight set;
23       $\mathcal{S}_{\lambda}^{j}\leftarrow Strategy(\mathcal{G}^{j-1},\delta_{p}^{j-1},x_{b},\mathcal{L}^{adv},\mathcal{S}_{\mathcal{M}})$;
24       // Updating the Parameters;
25       $\mathcal{G}^{j},\delta_{p}^{j}\leftarrow Update(\mathcal{G}^{j-1},\delta_{p}^{j-1},x_{b},\mathcal{L}^{adv},\mathcal{S}_{\lambda}^{j},\mathcal{S}_{\mathcal{M}})$;
26
27 end for
[…]
29 $\delta_{p}*\leftarrow\delta_{p}^{N}$;
Return $\mathcal{G}*,\delta_{p}*$
Algorithm 1 Training of the Perturbation Generator
III-E ID-Guard for Adversarial Training
A potential application of ID-Guard is leveraging the proposed perturbation generator $\mathcal{G}$ as an adversarial attack module for adversarial training of facial manipulation models to enhance their robustness. During adversarial training, the well-trained $\mathcal{G}$ can rapidly generate training images with adversarial patterns, which are then used to train the facial manipulation model $\mathcal{M}$. Meanwhile, $\mathcal{G}$ is simultaneously fine-tuned to adapt to $\mathcal{M}$.
Specifically, adversarial training can be formulated as a bi-level min-max optimization problem, where the inner maximization aims to generate adversarial examples that maximize the loss, while the outer minimization seeks to update the model parameters to minimize the bad-case loss. Given a facial manipulation model $\mathcal{M}$, it can be formulated as:
$\min_{{\theta}_{\mathcal{M}}}\mathbb{E}_{(x,a)}\left[\max_{{\theta}_{\mathcal{G}},\mathcal{G}(x)\in\mathcal{S}}\mathcal{L}(\mathcal{M}(x+\mathcal{G}(x),a))\right],$ (15)
where, the adversarial perturbation $\delta=\mathcal{G}(x)$ is constrained within a $\ell_{\infty}$-bounded perturbation set $\mathcal{S}$. The inner maximization problem optimizes the perturbation generator $\mathcal{G}$ to produce adversarial perturbations that maximize the loss. Specifically, the adversarial loss for the selected facial manipulation model $\mathcal{M}$ is computed in this step. The outer minimization then updates the parameters of $\mathcal{M}$ to enhance its robustness against such perturbations. Here, the training loss of $\mathcal{M}$ defined in the specific algorithm is computed. We will discuss its effectiveness in Section IV-D5.
@@ IV-A1 Datasets
In our experiments, the CelebAMask-HQ dataset [48] is selected to train the perturbation generator. It consists of more than 30,000 face images, where each image carries semantic masks for 19 facial component categories. These fine-grained mask labels can provide support for computing the binary mask loss during the training stage. To adequately evaluate the performance as well as the generalizability of our method and the competing algorithms, we test them on three datasets, including CelebAMask-HQ [48], LFW [49], and FFHQ [50].
Figure 5: Visual example of $L_{2}^{face}$ metric design. (a) is a natural image, (b) is a forged image, and (c) and (d) are disrupted images in two different situations.
@@ IV-A2 Target Models
Ours (S-II) 0.534 0.005 1.000 0.330 0.103 0.970 0.509 0.023 1.000 0.348 0.028 0.995 0.184 0.082 0.963
Ours (P-Pert) 0.213 0.052 1.000 0.301 0.090 0.978 0.455 0.046 1.000 0.063 0.246 0.637 0.118 0.262 0.730
IV-A3 Baselines
To demonstrate the superiority of the proposed method in face identity protection and cross-model transferable performance, four advanced proactive defense methods including Disrupting [12], PG [10], CMUA [11], and IAP [14], are selected as competing algorithms. Disrupting [12] disrupts facial manipulation by iteratively solving gradient-based adversarial perturbations on the target model. PG [10] achieves transferable adversarial perturbation generation in gray-box scenarios by attacking a surrogate model. CMUA [11] is a baseline of universal defense against multiple models. IAP [14] designs an information-containing adversarial perturbation. For a fair comparison, we implement only its proactive disruption component, while the information embedding and extraction aspects are discussed in Section IV-E.
IV-A4 Metrics
where $(i,j)$ is the coordinate of pixels and $Face_{i,j}$ is the binary facial mask of the image. The pixel value of its face area is $1$, otherwise, it is 00. The binary facial mask is calculated by Dlib  [^5] https://pypi.org/project/dlib. The intuition behind this design is that the traditional full-image $L_{2}$ distance metric fails to reflect distortions in the facial region. As illustrated in Fig. 5, although the disrupted image in (c) is reported as successful under the $L_{2}$ distance metric, the perturbation primarily affects the background while leaving the facial region largely intact. In contrast, the proposed $L_{2}^{face}$ metric provides a more accurate measure of identity-related distortions in the facial region, which aligns with human perception. Additionally, we evaluate the identity similarity (noted as ID sim. in Tables) computed by Arcface [46] between the forged and distorted outputs. Defense success rates are also considered. Previous works [12, 11] have generally determined the success of a defense by whether the $L_{2}$ distance is greater than $0.05$, but this is incomplete in the task of preventing face stigmatization. As shown in Fig. 6, the distorted output in Fig. 6 (c) reports a successful defense at the $L_{2}^{face}$ distance, but it seems to only blacken the face without destroying the individual’s identity. Therefore, we propose that both $L_{2}^{face}$ distance greater than $0.05$ and identity similarity less than $0.4$ be satisfied to indicate successful defense, which is a more challenging evaluation. The contrast between (d) and (e) in Fig. 6 shows the necessity of considering both restrictions simultaneously.
IV-A5 Implementation Details
Figure 7: Visual examples of disruption to different facial manipulations.
@@ IV-B Comparison with Baselines
Table I summarizes the quantitative comparison of the proposed ID-Guard with competitive algorithms. Our method is reported separately under two strategies, as presented in III-C. The Disrupting [12] produces gradient-based adversarial perturbations against StarGAN [1], which makes it overfit in disrupting this model at the cost of cross-model performance, as shown in Fig. 7. The perturbation optimization of CMUA [11] and PG [10] is unconstrained and thus has limited destruction to the identity semantics. The visual example of the CMUA in Fig. 7 shows that the person’s identity in the distorted image can still be recognized. Attributed to the feature correlation measurement loss employed, IAP [14] improves the performance to destroy identification to some extent. In comparison, our method significantly destroys the identity semantics of images and effectively prevents the stigmatization of faces. Furthermore, the baseline methods equally weight the attack loss of different facial manipulations, leading to perturbations biased towards vulnerable models. Due to the introduction of the dynamic parameter strategy, we achieve balanced performance against various facial manipulations. For the most robust model against attack, HiSD [5], ID-Guard improves the defense success rate by $50\%$ compared with the state-of-the-art method.
On the three selected datasets, ID-Guard under both strategies demonstrated superior performance. For strategy I, the MGDA algorithm is employed to adjust the weights, which has the advantage of eliminating the need for human intervention and prior knowledge during the training process. However, this non-intervention makes MGDA-based ID-Guard exhibit an extreme effect to some extent: it performs better on the more vulnerable model (e.g., AGGAN [2]) and the more robust model (e.g., HiSD [5]), but does not provide significant improvement on models in the middle (e.g., RelGAN [4]). Moreover, additional backpropagation computation is required at each iteration when implementing MGDA, which increases the training overhead. In contrast, the KPI-based strategy balances different models by maintaining a set of prior parameters, ensuring more stable defense performance across models and a more balanced performance improvement. However, its drawback is that preliminary experiments are required before formal training to determine the values of the prior weight set.
We conducted an additional evaluation of the optimized gradient-prior perturbation, P-Pert. Similar to CMUA, P-Pert is a cross-model universal adversarial perturbation designed to protect multiple facial images. As shown in the quantitative results Table I and the visualizations in Fig. 7, P-Pert significantly outperforms the baseline universal perturbation in both facial region disruption and cross-model balance. This advantage stems from its use of a dynamic weight strategy and an identity disruption module for optimization. Compared to the proposed generative perturbation, P-Pert achieves cross-image universality at the cost of a certain degree of reduced defense performance. Defenders can balance performance and computational efficiency by selecting either $\mathcal{G}$ or P-Pert for image protection.